OpenStudy (pottersheep):

PLEASE HELP!! How do I do this Grade 9 Triangle question???

OpenStudy (pottersheep):

I cant find e and c!

OpenStudy (anonymous):

can you take a better pic? or draw it please?

OpenStudy (pottersheep):

This is the only picture I have sent from a student I am supposed to tutor..sorry! :(

OpenStudy (anonymous):

oh that's fine

OpenStudy (anonymous):

I'll try my best to read it

OpenStudy (anonymous):

brb

OpenStudy (pottersheep):

thank you!

OpenStudy (anonymous):

b = 50 .. a = 180-2(50) = 80 b+d = 180 => d = 130 c+e = 180-130 = 50

OpenStudy (anonymous):

hey, is WX = UW ? can't see!

OpenStudy (pottersheep):

apparently not! :( but I think it has to be...

OpenStudy (anonymous):

Yep! If so, then you have c=e or c=e=50/2 = 25 ..

OpenStudy (anonymous):

This is what I have so far: b=50 a=180-50(2) a=80 180-50 d=30

OpenStudy (pottersheep):

Thank you guys :)